OpenStudy (anonymous):

How do I solve this absolute equation? |4f + 1| -2 = 5 The first answer, when the problem is in it's original state, is obviously f = 1.5 and I know the second answer needs to be 2, but how would I get to that so the answer is 2?

OpenStudy (zale101):

isolate the aboslute value from -2, that's by adding +2 to the right side

OpenStudy (zale101):

can u add 2 to both sides?

OpenStudy (anonymous):

yes, you need to add two to each side first. Once you have done that, you will need to solve for positive and negative values such that: (4f + 1)= 5+2 -(4f+1)= 5+2

OpenStudy (zale101):

right @derk765

OpenStudy (anonymous):

I solved the equation, thank for your inputs!

OpenStudy (anonymous):

Therefore, there should be two separate answers. You should be able to finish the rest using simple algebra and the distributive property

OpenStudy (anonymous):

Glad we could help!!